# Gated Rotary-Enhanced Linear Attention for Long-term Sequential Recommendation

### v1

@@ Abstract.
In Sequential Recommendation Systems (SRSs), Transformer models show remarkable performance but face computation cost challenges when modeling long-term user behavior sequences due to the quadratic complexity of the dot-product attention mechanism. By approximating the dot-product attention, linear attention provides an efficient option with linear complexity. However, existing linear attention methods face two limitations: 1) they often use learnable position encodings, which incur extra computational costs in long-term sequence scenarios, and 2) they may not consider the user’s fine-grained local preferences and confuse these with the actual change of long-term interests. To remedy these drawbacks, we propose a long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention (RecGRELA). Specifically, we first propose a Rotary-Enhanced Linear Attention (RELA) module to model long-range dependency within the user’s historical information using rotary position encodings. We then introduce a local short operation to incorporate local preferences and demonstrate the theoretical insight. We further introduce a SiLU-based Gated mechanism for RELA (GRELA) to help the model determine whether a user’s behavior indicates local interest or a genuine shift in long-term preferences. Experimental results on four public datasets demonstrate that our RecGRELA achieves state-of-the-art performance compared to existing SRSs while maintaining low memory overhead.
Sequential Recommendation, Linear Attention, Rotary Position Encoding, Gated Mechanism
ccs: Information systems Recommender systems
1. Introduction
Figure 1. Illustration of differences of SASRec, LinRec, and our RecGRELA. $\phi(\cdot)$ denotes the kernel function. (a) SASRec utilizes Softmax attention, where similarities between all query-key pairs are computed, leading to quadratic complexity for sequence length. (b) LinRec implements linear attention by transforming Q and K with $\phi(\cdot)$. This reordering of computations achieves linear complexity, though it may diminish the model’s expressive power. (c) Our RecGRELA enhances linear attention by incorporating Rotary Positional Encoding (RoPE) for Q and K and adding a shortcut connection, aiming to achieve strong expressive capability while maintaining linear computational complexity and removing the need for explicit Learned Position Encodings (LPE).
@@ 1. Introduction
Sequential Recommender Systems (SRSs) (Wang et al., 2019; Fang et al., 2020; Pan et al., 2024; Niu et al., 2025; Zhang et al., 2025; Yang et al., 2023; Zhang et al., 2021) have become a fundamental technology in personalized services, significantly transforming how online platforms predict user behavior. By analyzing sequences of user interactions across e-commerce, streaming services, and social media, SRSs capture both short-term interests and long-term preference trends. However, modeling long-term dependencies is challenging, as efficiently handling long sequences demands a balance between computational cost and model performance, while also capturing the dynamic and complex patterns in user interests. To address the challenge of modeling user behavior sequences, researchers explore various methodologies. Traditional approaches, such as those based on Markov Chains (He et al., 2017; Rendle et al., 2010) or Recurrent Neural Networks (RNNs) (Chen et al., 2018; Cui et al., 2018; Donkers et al., 2017; Xu et al., 2019), lay the groundwork but also face limitations in capturing long-range dependencies or suffered from computational inefficiencies during training and inference, particularly with increasing sequence lengths. Consequently, the field moves towards more advanced models that can handle complex sequential dependencies. This also presents several new challenges for the field.
Specifically, the first challenge is Efficiency. Transformer models (Kang and McAuley, 2018; Sun et al., 2019; Zhang et al., 2019) emerge as a compelling replacement for earlier approaches like RNNs, leveraging their ability to process sequential data and capture both local and global preferences of users. However, the dot-product attention (as show in Fig. 1(a)) in Transformers exhibits quadratic complexity relative to sequence length, resulting in high costs for long sequences. To address this issue, linear attention (Katharopoulos et al., 2020; Liu et al., 2023) has been proposed as an efficient alternative, providing linear complexity (as show in Fig. 1(b)). However, compared to dot-product attention, current linear attention approaches may suffer from performance drops.
@@
In addition to the efficiency of attention mechanisms, effectively capturing the positional information in user interaction sequences is another crucial challenge that affects both model performance and operational efficiency. Position encoding is widely adopted to capture the sequential order of items users interact with due to its ability to enhance the positional awareness of contextual vectors. As illustrated in Fig. 1(a) and (b), the SASRec (Kang and McAuley, 2018) and LinRec (Liu et al., 2023) both rely on learnable position encoding. The learned positional encoding requires more parameters than fixed positional encoding schemes since each position needs its own learned vector, and they do not have any inherent structure that captures the notion of relative distances between positions. To handle these drawbacks, some works (Su et al., 2024; Tian et al., 2024) employ an improved Rotary Position Encoding (RoPE), which injects relative positional information in a more efficient manner. However, they still use the dot-product attention mechanism, which limits its scalability for long sequences. Furthermore, the RoPE exhibits inefficiency in multi-head attention scenarios, restricting the model’s overall expressive capability.
@@
Linear attention, while effective for grasping overarching long-term dependencies, local preferences are equally vital for the long-term sequential recommendation, driven by users’ immediate preferences. For instance, a user might intensely engage with a specific series of related items. They could watch multiple episodes of The Avengers and then immediately consume supplementary content detailing Marvel’s production process. Subsequently, they might delve into documentaries that explore the creators behind such cinematic universes. However, linear attention’s tendency to smooth features can obscure important local information, while RoPE’s distance-based decay (Su et al., 2024) weakens attention to nearby items, particularly in long sequences. Both these characteristics hinder the precise modeling of local patterns. Even when linear attention successfully captures both long-term interests and short-term local preferences, a significant challenge persists: when a user’s preferences really change, it’s hard for the model to identify them. For instance, consider a user like historical documentaries over a weekend, binge-watches several interconnected Marvel movies like The Avengers series, and then explores some behind-the-scenes content about their production. A model that struggles to differentiate signals might misinterpret this intense but potentially short-lived burst of activity as a long-term shift in preference. Consequently, it might aggressively recommend a deluge of Marvel-related content for weeks – from other superhero films they’ve already seen, to comic book adaptations, or even tangentially related action movies. This happens because the model overweights fleeting interests. The user might simply want a weekend of superhero fun and is now looking for a historical drama, getting frustrated by the repetitive and misaligned suggestions.
@@
To solve the aforementioned challenges, we propose RecGRELA, a novel long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention block (as shown in Fig. 1(c)) to effectively capture long-range dependency. Specifically, we design a Rotary Enhanced Linear Attention (RELA) module. RELA builds upon linear attention, efficiently solving the computation limitation with a linear complexity of $\mathcal{O}(N)$, It also enhances its positional awareness by incorporating Rotary Position Encoding (RoPE). RoPE enables the model to capture relative positional relationships between items in a parameter-efficient manner in multi-head scenarios, overcoming the drawback of learnable positional encodings that introduce additional parameters. Concurrently, we design a Local Shortcut operation with causal 1D convolution in RELA to integrate local preferences and present theoretical insights based on the matrix rank analysis. We further introduce a SiLU-based gated mechanism to identify whether user behavior is a temporary burst of interest or a true shift of long-term interest via adjusting the output of RELA. With this gated mechanism, we remove the linear projection on the value matrix of the attention module to reduce the computational cost.
@@
  • We further develop a Local Shortcut operation to inject dense local interaction into the linear attention framework. This mechanism combines causal 1D convolution, explicitly modeling local preferences of user behaviors without sacrificing computational efficiency. We introduce a SiLU-based gated linear attention mechanism to distinguish true shifts in user preferences from transient local interests.
@@
  • Based on these designs, we further construct the RecGRELA, a novel long-term sequential recommendation paradigm that achieves competitive performance compared with state-of-the-art SRSs on four real-world datasets, keeping low memory overhead. Our source code is available at Github111 https://github.com/TBI805/RecGRELA..
Table 1. Notations and their descriptions for Section 2.
Notations Description
$\mathcal{U}$ user set with each user $u\in\mathcal{I}$
$\mathcal{V}$ user set with each user $v\in\mathcal{V}$
$\mathcal{S}_{u}$ interaction sequence for user $u$
$N$ sequence length
$D,d$ embedding dimension and head dimension of training sequence
$\boldsymbol{X}$ input sequence
$\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}$ projection matrix for $\boldsymbol{X}$
$\boldsymbol{Q}_{m},\boldsymbol{K}_{m},\boldsymbol{V}_{m}$ query, key and value matrix on position $m$
$\boldsymbol{O}_{m}$ attention result on position $m$
$\mathcal{O}\left(\cdot\right)$ computation complexity
$\phi\left(\cdot\right)$ kernel function
$\mathbf{R}$ the rotation matrix
$\mathcal{C}\left(\cdot\right)$, $\mathcal{C}^{-1}\left(\cdot\right)$ operator for converting between real and complex vector representations
$\mathcal{R}\left(\cdot\right)$ RoPE operation
$\mathrm{Re}[\cdot]$ real part of the complex vector
2. PRELIMINARY
2.1. Problem Definition
[…]
2.2. Dot-Product Attention and Linear Attention
@@
Let $\boldsymbol{X}\in\mathbb{R}^{N\times D}$ denote an input sequence. Formally, the input sequence $\boldsymbol{X}$ is projected by three projection matrices $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}\in\mathbb{R}^{D\times d}$, to corresponding representations $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}_{Q}$, $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}_{K}$, and $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}_{V}$. The dot-product attention (Vaswani et al., 2017) can be written as:
(1) $\quad{\boldsymbol{O}}_{m}=\sum_{n=1}^{N}\frac{\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}{\sum_{n=1}^{N}\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}\boldsymbol{V}_{n},$
@@
where $\boldsymbol{Q}_{m},\boldsymbol{K}_{n},\boldsymbol{V}_{n}\in\mathbb{R}^{1\times d}$, and $d$ is the head dimension. In this case, calculating attention weights in dot-product attention leads to the computation complexity of $\mathcal{O}(N^{2})$ with respect to the input sequence length $N$. For autoregressive modeling, dot-product attention can be implemented by incorporating a masking mechanism, which restricts the influence on the $m$-th position to a position $n$ satisfying $n\leq m$.
@@
Linear attention (Katharopoulos et al., 2020) is proposed as an effective alternative that reduces the computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ compared with dot-product attention. Specifically, linear attention modifies the dot-product attention by replacing the non-linear Softmax operation with linear normalization, while incorporating an additional kernel function $\phi$ in query matrix $\boldsymbol{Q}$ and key matrix $\boldsymbol{K}$:
(2) $\boldsymbol{O}_{m}=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}{\sum_{n=1}^{N}\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}\boldsymbol{V}_{n}.$
@@
(3) $\boldsymbol{O}_{m}=\frac{\phi\left(\boldsymbol{Q}_{m}\right)\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$
where the computational complexity is decreased to $\mathcal{O}(N)$. In (Katharopoulos et al., 2020), the kernel function employs the Exponential Linear Unit (ELU) activation (Clevert, 2015), which prevents gradient vanishing when processing negative values of $\boldsymbol{X}$. In our work, we follow this setting for a fair comparison.
2.3. Position Encoding
Absolute Position Encoding (APE) (Vaswani et al., 2017) incorporates fixed values into the embeddings of sequence elements, establishing a consistent reference frame across all tokens. For a token at the position $t$ within a sequence of length $N$, this encoding is computed through a deterministic function utilizing paired sine and cosine operations across different frequencies. The formulation is denoted as follows:
(4) $PE\left(p,2t\right)$ $=\sin\left(p/10000^{2t/\left(D/2\right)}\right),$
@@ 2.3. Position Encoding
Learnable Position Encoding (LPE) (Wang et al., 2021; Kang and McAuley, 2018) adopts a distinct methodology by modeling positional information as learnable parameters that are optimized during the training process. Specifically, for an item embedding x at position $m$, the transformation is performed as follows:
(5) $\textbf{{x}}_{m}=\textbf{{x}}+p_{m},$
@@
Rotary Position Encoding (RoPE) (Su et al., 2024) captures relative positional information (Shaw et al., 2018) through absolute position encoding methods. To achieve this, RoPE encodes the relative positional information via a rotation matrix $\mathbf{R}\in\mathbb{C}^{N\times(d/2)}$:
(6) $\mathbf{R}\left(m,t\right)=e^{i\theta_{t}m},$
[…]
(7) $\bar{\boldsymbol{Q}}^{\prime}_{m}$ $=\boldsymbol{\bar{Q}}_{m}\odot\mathbf{R}\left(m,t\right),$
@@
(8) $\mathcal{R}\left(\boldsymbol{X}\right)=\mathcal{C}^{-1}\left(\mathcal{C}\left(\boldsymbol{X}\right)\odot\mathbf{R}\right),$
@@
Table 2. Notations and their descriptions for Section 3.
Notations Description
$\boldsymbol{E}$ learnable embedding matrix
$\boldsymbol{H}_{o}$, $\boldsymbol{H}$ initial and final embedding matrix
$\boldsymbol{H}_{g}$ input of the the GRELA block
$\boldsymbol{H}_{L}$ final output for prediction
$\boldsymbol{O}_{f}$ output of the RELA
$\boldsymbol{S}$ output of the causal convolution
$\operatorname{CSC}(\cdot)$ causal convolution operation
$\boldsymbol{M}_{\operatorname{CSC}}$ sparse matrix associated with causal convolution and non-linear activation function
$\boldsymbol{M}_{eq}$ equivalent full attention map
$\boldsymbol{W}_{1}$, …, $\boldsymbol{W}_{5}$ weight matrix for linear projection
$\boldsymbol{b}_{1},...,\boldsymbol{b}_{5}$ bias for linear projection
$\boldsymbol{O}_{i}$ attention result on position $i$
$\hat{\mathcal{Y}}$ predicted recommendation result
$\mathcal{L}_{CE}$ cross-entropy loss
3. METHODOLOGY
In this section, we introduce our proposed framework, RecGRELA. The framework overview is shown in Figure 2.

### v2

# Methodology, Writing – original draft, Investigation, Software

@@ Abstract
In Sequential Recommendation Systems (SRSs), Transformer models have demonstrated remarkable performance but face computational and memory cost challenges, especially when modeling long-term user behavior sequences. Due to its quadratic complexity, the dot-product attention mechanism in Transformers becomes expensive for processing long sequences. By approximating the dot-product attention using elaborate mapping functions, linear attention provides a more efficient option with linear complexity. However, existing linear attention methods face three limitations: 1) they often use learnable position encodings, which incur extra computational costs in long-term sequence scenarios, 2) they may not sufficiently account for user’s fine-grained local preferences (short-lived burst of interest), and 3) they try to capture some temporary activities, but often confuse these with stable and long-term interests. This can result in unclear or less effective recommendations. To remedy these drawbacks, we propose a long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention (RecGRELA). Specifically, we first propose a Rotary-Enhanced Linear Attention (RELA) module to efficiently model long-range dependency within the user’s historical information using rotary position encodings. Then, we introduce a local short operation to add the local preferences of interactions and show the theoretical insight. We further introduce a SiLU-based Gated mechanism for RELA (GRELA) to help the model tell if a user behavior shows a short-term, local interest or a real change in their long-term tastes. Experimental results on four public benchmark datasets show that our RecGRELA achieves state-of-the-art performance compared with existing SRSs based on Recurrent Neural Networks, Transformer, and Mamba while keeping low memory overhead. Our source code is available at https://anonymous.4open.science/r/RecGLARE-B4AC.
keywords:
[…]
{highlights}
Rotary Enhanced Linear Attention integrates RoPE for efficient long-sequence modeling.
Local Shortcut operation injects dense local user preferences via causal convolution.
RecGRELA achieves competitive performance with low memory overhead.
1 Introduction
Figure 1: Illustration of differences of SASRec, LinRec, and our RecGRELA. $\phi$ denotes the kernel function.
@@ 1 Introduction
Sequential Recommender Systems (SRSs) [54, 9, 37, 35, 68, 48] have become a fundamental technology in personalized services, significantly transforming how online platforms predict user behavior. By analyzing sequences of user interactions across e-commerce, streaming services, and social media, SRSs capture both short-term interests and long-term preference trends. However, modeling long-term dependencies is challenging, as efficiently handling long sequences demands a balance between computational cost and model performance, while also capturing the dynamic and complex patterns in user interests.
To address the challenge of modeling user behavior sequences, researchers explore various methodologies. Traditional approaches, such as those based on Markov Chains [17, 42] or Recurrent Neural Networks (RNNs) [1, 3, 4, 58], lay the groundwork but also face limitations in capturing long-range dependencies or suffered from computational inefficiencies during training and inference, particularly with increasing sequence lengths. Consequently, the field moves towards more advanced models that can handle complex sequential dependencies more effectively. This also presents new challenges.
@@
Specifically, the first challenge is Efficiency. Transformer models [19, 47, 63] emerge as a compelling replacement for earlier approaches like RNNs, leveraging their ability to process sequential data and capture both local and global preferences of users. However, the dot-product attention mechanism in Transformers exhibits quadratic complexity relative to sequence length, resulting in high costs for long sequences. To address this issue, linear attention [20, 31] has been proposed as an efficient alternative, providing linear complexity while maintaining the modeling capabilities. Position encoding is widely adopted to capture the sequential order of items users interact with due to its ability to enhance the positional awareness of contextual vectors. As illustrated in Fig. 1(a) and (b), the above attention-based methods rely on learnable position encoding. The learned positional encoding requires more parameters than fixed positional encoding schemes since each position needs its own learned vector, and they do not have any inherent structure that captures the notion of relative distances between positions. To handle these drawbacks, some works [46, 50] employ an improved Rotary Position Encoding (RoPE), which injects relative positional information in a more efficient manner. However, they still use the dot-product attention mechanism, which limits its scalability for long sequences. Furthermore, the RoPE exhibits inefficiency in multi-head attention scenarios, restricting the model’s overall expressive capability.
Linear attention, while effective for grasping overarching long-term dependencies, can struggle to identify Fine-Grained Local Preferences. In recommendation contexts, user interactions often reveal dense, localized activity sequences. For instance, a user might intensely engage with a specific series of related items. They could watch multiple episodes of The Avengers and then immediately consume supplementary content detailing Marvel’s production process. Subsequently, they might delve into documentaries that explore the creators behind such cinematic universes. These closely related, short-term sequences of engagement are crucial for understanding immediate user intent, yet many existing approaches overlook their significance in long-term sequential recommendation.
@@
Even when linear attention successfully captures both long-term interests and short-term local preferences, a significant challenge persists: when a user’s preferences really change, it’s hard for the model to identify them. For instance, consider a user like historical documentaries over a weekend, binge-watches several interconnected Marvel movies like The Avengers series, and then explores some behind-the-scenes content about their production. A model that struggles to differentiate signals might misinterpret this intense but potentially short-lived burst of activity as a long-term shift in preference. Consequently, it might aggressively recommend a deluge of Marvel-related content for weeks – from other superhero films they’ve already seen, to comic book adaptations, or even tangentially related action movies. This happens because the model overweights fleeting interests. The user might simply want a weekend of superhero fun and is now looking for a historical drama, getting frustrated by the repetitive and misaligned suggestions.
@@
To solve the aforementioned challenges, we propose RecGRELA, a novel long-term sequential Recommendation model with Gated Rotary Enhanced Linear Attention block (as shown in Fig. 1(c)) to effectively capture long-range dependency. Specifically, we design a Rotary Enhanced Linear Attention (RELA) module. RELA builds upon the efficiency of linear attention, directly addressing the quadratic complexity issue of standard Transformers, but significantly enhances its positional awareness by incorporating Rotary Position Encoding (RoPE). This allows the model to effectively capture relative inter-item positional relationships in a parameter-efficient manner, overcoming the drawbacks of purely learnable or fixed encodings and the inefficiency of RoPE in standard multi-head dot-product attention. Concurrently, it integrates local preferences efficiently via a Local Shortcut operation with causal 1D convolution, and we present theoretical insights based on matrix rank analysis. We further introduce a SiLU-based gated mechanism to identify whether user behavior is a temporary burst of interest or a true shift of long-term interest via adjusting the output of RELA. With this gated mechanism, we remove the linear projection on the value matrix of the attention module to reduce the computational cost.
@@
  • We develop a Local Shortcut operation to inject dense local interaction into the linear attention framework. This mechanism combines causal 1D convolution, explicitly modeling local preferences of user behaviors without sacrificing computational efficiency.
@@
  • We introduce a SiLU-based gated linear attention mechanism to distinguish true shifts in user preferences from transient local interests. Based on these designs, we further construct the RecGRELA, a novel long-term sequential recommendation paradigm that achieves competitive performance compared with state-of-the-art SRSs on four benchmark datasets, keeping low memory overhead.
2 PRELIMINARY
2.1 Problem Definition
Considering a set of users $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$ and a set of items $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$. We use $\mathcal{S}_{u}=\left\{v_{t}\right\}_{t=1}^{T}$ to denote an interaction sequence of $T$ chronologically ordered items that the user $u$ has interacted with. Given a user interaction sequence $\mathcal{S}_{u}$, the task of sequential recommendation is to compute a ranking list consisting of Top-$k$ items that the user $u$ is likely to interact with at the next time step $T+1$. We define $N$ as the sequential length and $D$ as its corresponding dimension. Typically, the $N$ is often relatively short, but in some complex scenarios, where the $N$ satisfies $N$ > 1.5DD (e.g., when DD=64, the $N$ > 96), the sequence is regarded as “long-term[31].
2.2 Dot-Product Attention and Linear Attention
@@
Let $\boldsymbol{X}\in\mathbb{R}^{N\times D}$ denotes an input sequence. Formally, the input sequence $\boldsymbol{X}$ is projected by three projection matrices $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K},\boldsymbol{W}_{V}\in\mathbb{R}^{D\times d}$, to corresponding representations $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}_{Q}$, $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}_{K}$, and $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}_{V}$. The dot-product attention [52] can be written as:
$\quad{\boldsymbol{O}}_{m}=\sum_{n=1}^{N}\frac{\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}{\sum_{n=1}^{N}\exp\left(\boldsymbol{Q}_{m}\boldsymbol{K}_{n}^{\top}/\sqrt{d}\right)}\boldsymbol{V}_{n},$ (1)
@@
where $\boldsymbol{Q}_{m},\boldsymbol{K}_{n},\boldsymbol{V}_{n}\in\mathbb{R}^{1\times d}$, and $d$ is head dimension. In this case, calculating attention weights in dot-product attention leads to the computation complexity of $\mathcal{O}(N^{2})$ with respect to the input sequence length $N$. For autoregressive modeling, dot-product attention can be implemented by incorporating a masking mechanism, which restricts the influence on the $m$-th position to a position $n$ satisfying $n\leq m$.
@@
Linear attention [20] is proposed as an effective alternative that reduces the computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ compared with dot-product attention. Specifically, linear attention modifies the dot-product attention by replacing the non-linear Softmax operation with linear normalization, while incorporating an additional kernel function $\phi$ in query matrix $\boldsymbol{Q}$ and key matrix $\boldsymbol{K}$:
$\boldsymbol{O}_{m}=\sum_{n=1}^{N}\frac{\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}{\sum_{n=1}^{N}\phi\left(\boldsymbol{Q}_{m}\right)\phi\left(\boldsymbol{K}_{n}\right)^{\top}}\boldsymbol{V}_{n}.$ (2)
@@
$\boldsymbol{O}_{m}=\frac{\phi\left(\boldsymbol{Q}_{m}\right)\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\boldsymbol{V}_{n}\right)}{\phi(\boldsymbol{Q}_{m})\left(\sum_{n=1}^{N}\phi\left(\boldsymbol{K}_{n}\right)^{\top}\right)},$ (3)
where the computational complexity is decreased to $\mathcal{O}(N)$. In [20], the kernel function employs the Exponential Linear Unit (ELU) activation [2], which prevents gradient vanishing when processing negative values of $\boldsymbol{X}$. In our work, we follow this setting for a fair comparison.
2.3 Position Encoding
Absolute Position Encoding (APE) [52] incorporates fixed values into the embeddings of sequence elements, establishing a consistent reference frame across all tokens. For a token at the position $t$ within a sequence of length $N$, this encoding is computed through a deterministic function utilizing paired sine and cosine operations across different frequencies. The formulation is denoted as follows:
$PE\left(p,2t\right)$ $=\sin\left(p/10000^{2t/\left(D/2\right)}\right),$ (4)
@@ 2.3 Position Encoding
Learnable Position Encoding (LPE) [53, 19] adopts a distinct methodology by modeling positional information as learnable parameters that are optimized during the training process. Specifically, for an item embedding x at position $m$, the transformation is performed as follows:
$\textbf{{x}}_{m}=\textbf{{x}}+p_{m},$ (5)
@@
Rotary Position Encoding (RoPE) [46] captures relative positional information [43] through absolute position encoding methods via encoding the relative positional information with a rotation matrix $\mathbf{R}\in\mathbb{C}^{N\times(d/2)}$:
$\mathbf{R}\left(m,t\right)=e^{i\theta_{t}m},$ (6)
where $i$ denotes the imaginary unit in complex numbers. Following [52], the $\theta_{t}$ is set to $10000^{-t/(d/2)}$, where $t\in[1,2,...,d/2]$. Then RoPE converts $\boldsymbol{Q}_{m}$, $\boldsymbol{K}_{n}\in\mathbb{R}^{N\times{d}}$ to complex vector $\boldsymbol{\bar{Q}}_{m}$, $\boldsymbol{\bar{K}}_{n}\in\mathbb{C}^{N\times(d/2)}$ by considering (2tt)-th dim as real part and (2tt + 1)-th dim as imaginary part. The RoPE is applied to complex vectors of query and key as:
$\bar{\boldsymbol{Q}}^{\prime}_{m}$ $=\boldsymbol{\bar{Q}}_{m}\odot\mathbf{R}\left(m,t\right),$ (7)
@@
$\mathcal{R}\left(\boldsymbol{X}\right)=\mathcal{C}^{-1}\left(\mathcal{C}\left(\boldsymbol{X}\right)\odot\mathbf{R}\right),$ (8)
@@
3 METHODOLOGY